OpenStudy (anonymous):

Expand the following logarithm as much as possible: log[b]4x^2y

OpenStudy (anonymous):

2y(log[b]4+log[b]x) I'm not quite sure though...

OpenStudy (anonymous):

\[\log(4x^2y)=\log(4)+\log(x^2)+\log(y)=\log(4)+2\log(x)+\log(y)\]

OpenStudy (angela210793):